OpenStudy (imterribleatmath):

PLEASE HELP, Find the missing factor. 6x^2+13x-5=(2x+5)( ) blank is the missing factor.

OpenStudy (3mar):

May I help?

OpenStudy (3mar):

@imterribleatmath

OpenStudy (imterribleatmath):

yes please

OpenStudy (3mar):

Ok

OpenStudy (3mar):

\[6x^2+13x-5=(2x+5)( ... )\]

OpenStudy (3mar):

what is the first term in the second parentheses? for sure it (...)x^2

OpenStudy (3mar):

@imterribleatmath Can you kindly respond more faster, please?

OpenStudy (imterribleatmath):

what do u mean what is the second term i put it ?

OpenStudy (3mar):

6x^2=2x*??

OpenStudy (imterribleatmath):

the term in the second parentheses is the missing factor

OpenStudy (3mar):

yes, I know we are going to find it out! 6x^2=2x*??

OpenStudy (imterribleatmath):

honestly i have no clue where to start.

OpenStudy (3mar):

6x^2=2x*3x is not it?

OpenStudy (imterribleatmath):

i do not understand is that the missing factor?

OpenStudy (3mar):

\[3x*2x=6x^2\]

OpenStudy (3mar):

\[5=5*1\]

OpenStudy (imterribleatmath):

i'm sorry like i said, i'm clueless

OpenStudy (3mar):

Why you say that? It is simple. Trust me!

OpenStudy (imterribleatmath):

So could you give me then answer and kinda a explanation?

OpenStudy (3mar):

\[6x^2+13x-5=(2x+5)( 3x .. 1)\] can you choose + or - between 3x and 1?

OpenStudy (3mar):

@imterribleatmath

OpenStudy (imterribleatmath):

i can choose both

OpenStudy (3mar):